OpenStudy (anonymous):

It takes Emma 18 minutes to type and spell check 10 pages of a manuscript. Find how long it takes her to type and spell check 65 pages. a. 36 min c. 18 min b. 1170 min d. 117 min

OpenStudy (anonymous):

solve \[\frac{18}{10}=\frac{x}{65}\] is one way i.e. \[\frac{time}{page}=\frac{time}{page}\]

OpenStudy (anonymous):

another way is to say that 18 pages in 10 minutes is \(18\div 10=1.8\) pages per minute then \(65\times 1.8\) is your answer

OpenStudy (nirmalnema):

it takes 18 min for 10 pages.. so for 1 page....=18/10 and for 65 pages 18/10 *65 9*13=117 mins

OpenStudy (anonymous):

D

OpenStudy (anonymous):

yeah D